OpenStudy (anonymous):

The lengths of a triangle are 6,8,and10 can the triangle be a right triangle?

OpenStudy (agent0smith):

Check it using pythagoras theorem: \[c^2 = a^2+b^2\] where c is the hypotenuse (longest side) Check that \[10^2 = 6^2 + 8^2\]

OpenStudy (agent0smith):

If it satisfies that criteria, then it must be a right triangle.